# DeNOTS: Stable Deep Neural ODEs for Time Series

### v4

@@ 4.2.1 Error bounds for specific covariance functions
For some $\xi>0,Q>0$, consider the spectral density  [^3] At the limit $\xi\rightarrow 0$, the mean of the resulting GP tends to a natural cubic spline Golubev & Krymova (2013). The parameter $Q$ here represents our a priori beliefs about the data variance.: $F_{i}(\omega)=\frac{Q}{\omega^{4}+\xi^{4}}.$ Then, under Assumption 4.6, the interval error $\sigma_{\textrm{INT}}^{2}(\xi)$ from (8) has the following asymptotic:
$\frac{4\pi^{4}\sqrt{3}}{63}uQ\delta_{\min}^{4}\leq\underset{\xi\rightarrow 0}{\lim}\sigma_{\textrm{INT}}^{2}(\xi)\leq\frac{4\pi^{4}\sqrt{3}}{63}uQ\delta_{\max}^{4}.$
@@ 4.3 Long-term memory
On the other hand, forgetfulness could make the model more robust: time series often evolve quickly, so forgetting irrelevant observations may be beneficial Cao et al. (2019). To gauge this effect, we measure our method’s performance on perturbed sequences, with a small fraction replaced by standard Gaussian noise. The results are provided in Table 7 (row 2). Indeed, all the possibly forgetful models, i.e., the SNCDE versions and the GRU, are in the lead. Adding NF boosts robustness: Sync-NF and Anti-NF significantly outperform No-NF. Sync-NF performs slightly worse than Anti-NF. We hypothesise this is because the attack is uniformly distributed across the sequence: the corrupted tokens may appear closer to the end, in which case the Sync-NF model performs worse due to forgetfulness.
@@ 5 Main Experiments
Backbone UWGL InsectSound Pendulum Sepsis
GRU $0.5\pm 0.1$ $0.3\pm 0.2$ $0.73\pm 0.03$ $0.838\pm 0.004$
RoFormer $0.74\pm 0.04$ $0.29\pm 0.02$ $0.61\pm 0.02$ $0.924\pm 0.004$
Mamba $0.71\pm 0.07$ $0.41\pm 0.03$ $0.61\pm 0.03$ $0.829\pm 0.004$
GRU-ODE $0.4\pm 0.3$ $0.18\pm 0.03$ $0.6\pm 0.05$ $0.925\pm 0.003$
@@ 6 Conclusions
Our code is available at https://anonymous.4open.science/r/denots_neurips2025-7DE6/. It contains automated scripts to download all our datasets, and uses popular frameworks to make it recognisable to a wide audience of ML researchers, reproducing all our baselines in these frameworks. A detailed description of the datasets can be found in Appendix C.1. The baselines are described in Appendix C.2. More information about our pipeline can be found in Appendix C.3.
@@ Appendix A Related works
Neural ODEs provide an elegant way to take the sequence irregularity into account Oh et al. (2025). The representation evolves simultaneously with the observed system Chen et al. (2018). For example, GRU-ODE-Bayes De Brouwer et al. (2019) used Bayesian methods to account for missing observations optimally. To constrain the hidden trajectory to [1,1][-1,1], its authors introduced a strict NF, subtracting the current hidden state from the derivative.
As an improvement, the authors of Kidger et al. (2020) proposed to combine Neural ODEs with the theory of CDEs Friz & Victoir (2010) for time series classification. Instead of embedding the input data into the starting point, Neural CDEs interpolate it, using the resulting function to guide the trajectory. On the other hand, the original Neural CDEs from Kidger et al. (2020) are not sensitive to the irregularity of time intervals. So, the authors append the time difference as a separate feature. Moreover, these models have high memory usage. The controlling NN has to generate a matrix, which implies a weight dimension of $\sim d^{3}$, where $d$ is the dimensionality of the input features, instead of $d^{2}$ for conventional models, such as RNNs. DeNOTS employs a GRU cell as its dynamics, making it more memory efficient than Neural CDE, and sensitive to input time intervals, which is crucial for scaling time.
@@
Neural ODE stability is a well-researched topic. Prior work has approached it through adding specialized loss terms Kang et al. (2021); Rodriguez et al. (2022), constraining the form of the vector field Massaroli et al. (2020) or the eigenvalues of the NN weights Tuor et al. (2020). However, very little work has been done for Neural CDEs, since the problem is significantly complicated by the presence of an external ”control signal“. Most notably, Morrill et al. (2021b) processes the input sequence in a windowed fashion, aggregating each window into a certain set of features to reduce length, while Morrill et al. (2021a) highlights the importance of selecting a continuous interpolant for the well-posedness of the initial-value problem. We prove that our CDE-type model is input-to-state stable, providing a significantly more general and formal result.
@@
Our theoretical analysis investigates discretisation uncertainty, which manifests as GP variance. This involves estimating the squared error of an integral over a linear function of a GP with multiple outputs. For a specific assumption, we have a reasonable estimate of the quadratic risk that follows from Golubev & Krymova (2013). While alternatives exist Stein (2012); van der Vaart & van Zanten (2008); Zaytsev et al. (2018), they are unsuitable for our setting, with an additional ODE layer on top.
@@ B.1 Time scaling
Our model $F$ maps the continuous input data trajectory $\hat{\mathbf{x}}(\cdot):[0,T]\rightarrow\mathbb{R}^{u}$ to a hidden state $\mathbf{h}\triangleq\mathbf{h}(T)\in\mathbb{R}^{v}$, i.e. $F:C([0,T],\mathbb{R}^{u})\rightarrow\mathbb{R}^{v}$. We describe the class of functions $F$ in terms of Lipschitz constants, where $F$ is $L_{F}$-Lipschitz, if:
$\|F(\hat{\mathbf{x}}_{1})-F(\hat{\mathbf{x}}_{2})\|_{2}\leq\rho(\hat{\mathbf{x}}_{1},\hat{\mathbf{x}}_{2})L_{F},\;\;\;\textrm{where}\;\rho(\hat{\mathbf{x}}_{1},\hat{\mathbf{x}}_{2})\triangleq\int_{0}^{T}\|\hat{\mathbf{x}}_{1}(t)-\hat{\mathbf{x}}_{2}(t)\|_{2}dt.$ (2)
In this section, for simplicity, we assume that $\mathbf{h}(0)=\mathbf{0}_{v}$. To broaden the class of functions the model may represent, we must increase $L_{F}$. We can explicitly bound $L_{F}$, using $T$ and the Lipschitz constants of $\mathbf{g}_{\theta}$:
@@
Suppose, that the function $\mathbf{g}_{\theta}$ is Lipschitz w.r.t. $\hat{\mathbf{x}},\mathbf{h}$ with constants $M_{x},M_{h}$ respectively. Then, $F$ is $L_{F}$-Lipschitz in the sense of (2), where $L_{F}=\frac{M_{x}}{M_{h}}(e^{M_{h}T}-1)$:
$\|F(\hat{\mathbf{x}}_{1})-F(\hat{\mathbf{x}}_{2})\|_{2}\leq\rho(\hat{\mathbf{x}}_{1},\hat{\mathbf{x}}_{2})\frac{M_{x}}{M_{h}}(e^{M_{h}T}-1).$
@@
Suppose, $\mathbf{h}_{1},\mathbf{h}_{2}$ are the trajectories, corresponding to $\hat{\mathbf{x}}_{1},\hat{\mathbf{x}}_{2}$, respectively:
$\frac{d\mathbf{h}_{i}}{dt}=\mathbf{g}_{\theta}(\hat{\mathbf{x}}_{i}(t),\mathbf{h}_{i}(t)).$
We introduce the following notation for convenience:
$\Delta(t)\triangleq\mathbf{h}_{1}(t)-\mathbf{h}_{2}(t).$
Then, we can write the following differential equation for $\Delta$:
$\frac{d\Delta}{dt}=\mathbf{g}_{\theta}(\hat{\mathbf{x}}_{1}(t),\mathbf{h}_{1}(t))-\mathbf{g}_{\theta}(\hat{\mathbf{x}}_{2}(t),\mathbf{h}_{2}(t)).$ (11)
@@
The absolute derivative of the norm is always lower or equal to the norm of the derivative, according to the Cauchy-Schwartz inequality:

### v5

@@ 4.2.1 Interpolation error
For some $\xi>0,Q>0$, say that $\hat{\mathbf{x}}(t)$ is a Gaussian Process with the spectral density  [^3] At the limit $\xi\rightarrow 0$, the mean of the resulting GP tends to a natural cubic spline golubev2013interpolation. The parameter $Q$ here represents our a priori beliefs about the data variance.: $F_{i}(\omega)=\frac{Q}{\omega^{4}+\xi^{4}}.$ Also, assume that the considered sequence $S$ is infinite, and the GP covariance function is such that increasing interval sizes increases the expectation of the interval error from (11). Then, the interval error $\mathbb{E}\sigma_{\textrm{INT}}^{2}(\xi)$ has the following asymptotic:
$\frac{4\pi^{4}\sqrt{3}}{63}uQ\delta_{\min}^{4}\leq\underset{\xi\rightarrow 0}{\lim}\mathbb{E}\sigma_{\textrm{INT}}^{2}(\xi)\leq\frac{4\pi^{4}\sqrt{3}}{63}uQ\delta_{\max}^{4}.$
A full version of the reasoning behind the above Lemma is provided in the Appendix.
@@
To test the effect of interpolation error on our models, we perform drop attacks (dropping a fraction of input tokens and replacing them with NaNs) on the Pendulum dataset. According to Table 7 (first row), all SNCDEs are robust to drop attacks, with NF versions outperforming No NF. The Sync-NF model is slightly better than Anti-NF, likely due to a larger $b-aL_{h}$, but that difference is insignificant compared to other baselines.
@@ 4.3 Long-term memory
On the other hand, forgetfulness could make the model more robust: time series often evolve quickly, so forgetting irrelevant observations may be beneficial cao2019online. To gauge this effect, we measure our method’s performance on perturbed sequences, with a small fraction replaced by standard Gaussian noise. The results are provided in Table 7 (row 2). Indeed, all the possibly forgetful models, i.e., the SNCDE versions and the GRU, are in the lead. Adding NF boosts robustness: Sync-NF and Anti-NF significantly outperform No-NF. Sync-NF performs slightly worse than Anti-NF. We hypothesise this is because the attack is uniformly distributed across the sequence: the corrupted tokens may appear closer to the end, in which case the Sync-NF model performs worse due to forgetfulness.
@@ 5 Main Experiments
Table 9: Table with performance ranks (lower is better), built on our results. The rank of each backbone is one plus the number of other backbones, the score of which is better according to a one-sided T-test on the corresponding mean and deviations, with the p-value cutoff set to 0.1. Therefore, lower is better, a rank of 1.0 means that there are no better baselines, and the one in question is the best.
Backbone UWGL InsectSound Pendulum Sepsis Mean
DeNOTS (ours) 1 1 1 1 1.0
Sync-NF SNCDE (ours) 1 1 2 1 1.25
TempFormer 4 1 6 5 4.0
Neural CDE 1 8 2 5 4.0
Neural RDE 1 6 1 9 4.25
RoFormer 5 5 6 3 4.75
GRU 8 1 4 7 5.0
Mamba 6 1 6 8 5.25
GRU-ODE 8 7 6 3 6.0
@@ 6 Conclusions
Our code is available at https://anonymous.4open.science/r/denots_iclr2026-400E/. It contains automated scripts to download all our datasets, and uses popular frameworks to make it recognisable to a wide audience of ML researchers, reproducing all our baselines in these frameworks. A detailed description of the datasets can be found in Appendix C.1. The baselines are described in Appendix C.2. More information about our pipeline can be found in Appendix C.3.
@@ Appendix A Related works
Neural ODEs provide an elegant way to take the sequence irregularity into account oh2025comprehensive. The representation evolves simultaneously with the observed system chen2018neural. For example, GRU-ODE-Bayes de2019gru used Bayesian methods to account for missing observations optimally. To constrain the hidden trajectory to [1,1][-1,1], its authors introduced a strict NF, subtracting the current hidden state from the derivative.
As an improvement, the authors of kidger2020neural proposed to combine Neural ODEs with the theory of CDEs friz2010multidimensional for time series classification. Instead of embedding the input data into the starting point, Neural CDEs interpolate it, using the resulting function to guide the trajectory. On the other hand, the original Neural CDEs from kidger2020neural are not sensitive to the irregularity of time intervals. So, the authors append the time difference as a separate feature. Moreover, these models have high memory usage. The controlling NN has to generate a matrix, which implies a weight dimension of $\sim d^{3}$, where $d$ is the dimensionality of the input features, instead of $d^{2}$ for conventional models, such as RNNs. DeNOTS employs a GRU cell as its dynamics, making it more memory efficient than Neural CDE, and sensitive to input time intervals, which is crucial for scaling time.
@@
Neural ODE stability is a well-researched topic. Prior work has approached it through adding specialized loss terms kang2021stable; rodriguez2022lyanet, constraining the form of the vector field massaroli2020stable or the eigenvalues of the NN weights tuor2020constrained. However, very little work has been done for Neural CDEs, since the problem is significantly complicated by the presence of an external ”control signal“. Most notably, morrill2021neural processes the input sequence in a windowed fashion, aggregating each window into a certain set of features to reduce length, while morrill2021neuralonline highlights the importance of selecting a continuous interpolant for the well-posedness of the initial-value problem. We prove that our CDE-type model is input-to-state stable, providing a significantly more general and formal result.
Time.
To escape the instability, introduced by possibly large time frames, most prior works enforce time to be roughly [0, 1] by setting it directly chen2018neural; rubanova2019latent; dupont2019augmented, re-parametrizing it chen2020neural or predicting it within tight bounds massaroli2020dissecting. Others devise models invariant to time transforms kidger2020neural. To the best of our knowledge, all existing works do not consider time to be an important hyperparameter.
@@
Our theoretical analysis investigates discretisation uncertainty, which manifests as GP variance. This involves estimating the squared error of an integral over a linear function of a GP with multiple outputs. For a specific assumption, we have a reasonable estimate of the quadratic risk that follows from golubev2013interpolation. While alternatives exist stein2012interpolation; van2008rates; zaytsev2018interpolation, they are unsuitable for our setting, with an additional ODE layer on top.
@@ B.1 Time scaling
As mentioned in the introduction, we argue that scaling time benefits expressiveness. By expressiveness, we mean the broadness of the class of functions that our network can represent. Now, we introduce several definitions.
Differentially definable mappings from trajectories $\mathcal{F}$.
We say that $F_{\mathbf{g}}\in\mathcal{F}$, if there exists an ODE with the vector field $\mathbf{g}(\mathbf{x}(t),\mathbf{h}(t)):\mathbb{R}^{u}\times\mathbb{R}^{v}\rightarrow\mathbb{R}^{v}$, with the starting condition $\mathbf{h}(0)=\mathbf{h}_{0}$ and the solution $\mathbf{h}(t):\mathbb{R}\rightarrow\mathbb{R}^{v}$: $\frac{d}{dt}\mathbf{h}(t)=\mathbf{g}(\mathbf{h}(t),\mathbf{x}(t)),$ such that $F_{\mathbf{g}}(\mathbf{x}(\cdot),\mathbf{h}_{0};t)=\mathbf{h}(t)$. Not all functions on trajectories can be represented this way.
Start-corrected differentially-definable mappings,
i.e. the elements of $\mathcal{F}$, minus their initial state $\mathbf{h}_{0}$: $\mathring{\mathcal{F}}=\{\mathring{F}_{\mathbf{g}}(\mathbf{x}(\cdot),\mathbf{h}_{0};t)=F_{\mathbf{g}}(\mathbf{x}(\cdot),\mathbf{h}_{0};t)-\mathbf{h}_{0}|F_{\mathbf{g}}\in\mathcal{F}\}.$ Since we set $\mathbf{h}_{0}=0$ in most practical cases, the distinction is rather technical. This is mostly to account for Lipschitzness w.r.t. $\mathbf{h}$, as we will see below.
Lipschitz constraints.
We introduce two ways to constrain this class with Lipschitz constants: with $\mathbf{g}$ being $M_{x},M_{h}$-Lipschitz
$\mathring{\mathcal{F}}_{g}(M_{x},M_{h})=\{\mathring{F}_{\mathbf{g}}\in\mathring{\mathcal{F}}|\mathbf{g}-M_{x},M_{h}\text{-lipschitz w.r.t. }\mathbf{x},\mathbf{h}\},$
and with $\mathring{F}_{\mathbf{g}}(\ldots;t)$ being $L_{x}(t),L_{h}(t)$-lipschitz, where $L_{x}(t),L_{h}(t):\mathbb{R}\rightarrow\mathbb{R}_{+}$:
$\mathring{\mathcal{F}}_{F}(L_{x}(\cdot),L_{h}(\cdot))=\{\mathring{F}_{\mathbf{g}}\in\mathring{\mathcal{F}}|\mathring{F}_{\mathbf{g}}-L_{x}(t),L_{h}(t)\text{-lipschitz w.r.t. }\mathbf{x},\mathbf{h}_{0}\}.$
Note, that $\mathring{\mathcal{F}}_{g}(M_{x},M_{h})$ represents all mappings that our ODE can learn, given a Lipschitz-constrained Neural Network $\mathbf{g}$, while $\mathring{\mathcal{F}}_{F}(L_{x}(\cdot),L_{h}(\cdot))$ represents all the possible Lipschitz-mappings that we may want to learn.
Equipped with the above definitions, we can formulate the following:
@@
The classes $\mathring{\mathcal{F}}_{g}(M_{x},M_{h})$ and $\mathring{\mathcal{F}}_{F}(L_{x}(\cdot),L_{h}(\cdot))$ are equal, given the following relations between their arguments:
$L_{x}(t)=M_{x}\sqrt{\frac{1}{2M_{h}}\left(e^{2M_{h}t}-1\right)};L_{h}(t)=e^{M_{h}t}-1.$
@@
We conduct the proof in two steps: first proving the subset relation $\mathring{\mathcal{F}}_{g}(M_{x},M_{h})\subseteq\mathring{\mathcal{F}}_{F}(L_{x}(\cdot),L_{h}(\cdot))$, and then the converse $\mathring{\mathcal{F}}_{F}(L_{x}(\cdot),L_{h}(\cdot))\subseteq\mathring{\mathcal{F}}_{g}(M_{x},M_{h})$. Together, these relations will imply equality of the two considered classes.
Subset relation.
First, we prove the relation $\mathring{\mathcal{F}}_{g}(M_{x},M_{h})\subseteq\mathring{\mathcal{F}}_{F}(L_{x}(\cdot),L_{h}(\cdot))$. Let’s consider an $M_{x},M_{h}$-Lipschitz $\mathbf{g}$, and prove that the corresponding $\mathring{F}$ is $L_{x},L_{h}$-Lipschitz.
Say, we have two trajectories, $\mathbf{x}_{1}(t),\mathbf{x}_{2}(t)$, each producing their repsective $\mathbf{h}_{1,2}(t)$, with $\mathbf{h}_{1}(0)\neq\mathbf{h}_{2}(0)$. The distance between the corresponding $\mathring{F}$-s will be given by:
$\|(F(\mathbf{x}_{1},\mathbf{h}_{0},t)-F(\mathbf{x}_{1},\mathbf{h}_{0},0))-(F(\mathbf{x}_{2},\mathbf{h}_{0},t)-F(\mathbf{x}_{2},\mathbf{h}_{0},0))\|\triangleq\|\Delta(t)\|,$
@@
where we denote the expression under the norm by $\Delta$ for brevity. The absolute derivative of the norm is always less or equal to the norm of the derivative, which can be proven using the Cauchy-Schwartz inequality:
$\frac{d}{dt}\|\Delta\|_{2}\leq\left|\frac{d}{dt}\|\Delta\|_{2}\right|=\left|\frac{d}{dt}\sqrt{\Delta\cdot\Delta}\right|=\frac{\Delta\cdot\Delta^{\prime}}{\|\Delta\|_{2}}\leq\frac{\|\Delta\|_{2}\cdot\|\Delta^{\prime}\|_{2}}{\|\Delta\|_{2}}=\|\Delta^{\prime}\|_{2}=$
Here $\Delta^{\prime}$ is the derivative of $\Delta$ w.r.t. $t$. We can write it using $\mathbf{g}$, the starting points cancel out:
$\|\Delta^{\prime}(t)\|_{2}=\|\mathbf{g}(\mathbf{x}_{1}(t),\mathbf{h}_{1}(t))-\mathbf{g}(\mathbf{x}_{2}(t),\mathbf{h}_{2}(t))\|_{2}=$
We add and subtract a ”mixed“ term $\mathbf{g}(\mathbf{x}_{1}(t),\mathbf{h}_{2}(t))$:
$\|\mathbf{g}(\mathbf{x}_{1}(t),\mathbf{h}_{1}(t))-\mathbf{g}(\mathbf{x}_{1}(t),\mathbf{h}_{2}(t))+\mathbf{g}(\mathbf{x}_{1}(t),\mathbf{h}_{2}(t))-\mathbf{g}(\mathbf{x}_{2}(t),\mathbf{h}_{2}(t))\|_{2}\leq$
The norm of that sum may be bounded using the triangle rule:
$\leq\|\mathbf{g}(\mathbf{x}_{1}(t),\mathbf{h}_{1}(t))-\mathbf{g}(\mathbf{x}_{1}(t),\mathbf{h}_{2}(t))\|_{2}+\|\mathbf{g}(\mathbf{x}_{1}(t),\mathbf{h}_{2}(t))-\mathbf{g}(\mathbf{x}_{2}(t),\mathbf{h}_{2}(t))\|_{2}\leq$
And both of these two norm-terms can now be bounded by the Lipschitz property:
$\leq M_{x}\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{2}+M_{h}\|\mathbf{h}_{1}-\mathbf{h}_{2}\|_{2}=$
We add and subtract $\mathbf{h}_{1}(0)-\mathbf{h}_{2}(0)\triangleq\Delta_{0}$ in the second term:
$M_{x}\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{2}+M_{h}\|\mathbf{h}_{1}(t)-\mathbf{h}_{2}(t)-(\mathbf{h}_{1}(0)-\mathbf{h}_{2}(0))+(\mathbf{h}_{1}(0)-\mathbf{h}_{2}(0))\|_{2}\leq$
We now notice that we can substitute $\Delta$ in the second term, and upper-bound it with the triangle inequality
$\leq M_{x}\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{2}+M_{h}\|\Delta\|_{2}+M_{h}\|\Delta_{0}\|_{2},$
where we denote $\Delta_{0}\triangleq\mathbf{h}_{1}(0)-\mathbf{h}_{2}(0)$.
Putting the above reasoning together, we are left with the differential inequality:
$\frac{d}{dt}\|\Delta\|\leq M_{x}\|\mathbf{x}_{1}-\mathbf{x}_{2}\|+M_{h}\|\Delta\|+M_{h}\|\Delta_{0}\|.$
Using the standard ODE trick, we assume that $\Delta(t)=C(t)e^{M_{h}t}$, and substitute this into the inequality:
$\frac{d}{dt}\|\Delta\|=\frac{dC}{dt}e^{M_{h}t}+C(t)M_{h}e^{M_{h}t}\leq M_{x}\|\mathbf{x}_{1}-\mathbf{x}_{2}\|+M_{h}C(t)e^{M_{h}t}+M_{h}\|\Delta_{0}\|.$
The terms $M_{h}C(t)e^{M_{h}t}$ cancel out:
$\frac{dC}{dt}e^{M_{h}t}\leq M_{x}\|\mathbf{x}_{1}-\mathbf{x}_{2}\|+M_{h}\|\Delta_{0}\|.$
Now, we separate the variables, and integrate from $\tau=0$ to $t$ :
$\int_{\tau=0}^{t}dC=C(t)-C(0)\leq\left(M_{x}\int_{\tau=0}^{t}\|\mathbf{x}_{1}(\tau)-\mathbf{x}_{2}(\tau)\|e^{-M_{h}\tau}d\tau\right)+$
$M_{h}\|\Delta_{0}\|\int_{\tau=0}^{t}e^{-M_{h}\tau}d\tau.$
The integral without $\mathbf{x}_{1,2}(t)$ can be integrated analytically:
$\int_{\tau=0}^{t}e^{-M_{h}\tau}d\tau=\frac{1}{M_{h}}\left(1-e^{-M_{h}t}\right).$
The integral with $\mathbf{x}_{1,2}$ can be bounded using Cauchy-Schwartz:
$\int_{\tau=0}^{t}\|\mathbf{x}_{1}(\tau)-\mathbf{x}_{2}(\tau)\|e^{-M_{h}\tau}d\tau\leq$
$\leq\sqrt{\int_{\tau=0}^{t}\|\mathbf{x}_{1}(\tau)-\mathbf{x}_{2}(\tau)\|^{2}d\tau}\sqrt{\int_{\tau=0}^{t}e^{-2M_{h}\tau}d\tau}$
The first term is the $L_{2}$-distance we introduced earlier, $\rho$, the second one can be integrated analytically:
$\int_{\tau=0}^{t}e^{-2M_{h}\tau}d\tau=\frac{1}{2M_{h}}\left(1-e^{-2M_{h}t}\right).$
As a starting condition, we know that $\Delta(0)=0$ (since it is start-corrected), so $C(0)=0$. Now we have an inequality for $C(t)$:
$C(t)\leq\|\Delta_{0}\|_{2}\left(1-e^{-M_{h}t}\right)+M_{x}\rho(\mathbf{x}_{1},\mathbf{x}_{2})\sqrt{\frac{1}{2M_{h}}\left(1-e^{-2M_{h}t}\right)}.$
Multiplying it by $e^{M_{h}t}$ gives us the inequality for $\Delta(t)$:
$\|\Delta(t)\|\leq\|\Delta_{0}\|_{2}\left(e^{M_{h}t}-1\right)+M_{x}\rho(\mathbf{x}_{1},\mathbf{x}_{2})\sqrt{\frac{1}{2M_{h}}\left(e^{2M_{h}t}-1\right)}.$
To get the individual expression for $L_{x}$, we simply set $\Delta_{0}$ to zero. To get the expression for $L_{h}$, we set $\rho$ to zero.
Superset relation.
Now we prove the converse relationship: $\mathring{\mathcal{F}}_{F}(L_{x}(\cdot),L_{h}(\cdot))\subseteq\mathring{\mathcal{F}}_{g}(M_{x},M_{h})$.
Assume that we have $F\in\mathring{\mathcal{F}}_{F}(L_{x}(\cdot),L_{h}(\cdot))$, where for some $M_{x},M_{h}$:
$L_{x}(t)=M_{x}\sqrt{\frac{1}{2M_{h}}\left(e^{2M_{h}t}-1\right)};L_{h}(t)=e^{M_{h}t}-1.$
We need to prove that the corresponding $\mathbf{g}$ is $M_{x},M_{h}$-Lipschitz.
We start with the $x$ component.
Consider two input trajectories $\mathbf{x}_{1}\neq\mathbf{x}_{2}$, and equal starting points $\mathbf{h}_{1}=\mathbf{h}_{2}=\mathbf{h}_{0}$. Since $F$ is Lipschitz, we can write the following inequality: